OpenStudy (anonymous):

A 945 kg drag race car accelerates from rest to 117 km/h in 0.977 s. What change in momentum does the force produce? Answer in units of kg · m/s

OpenStudy (anonymous):

first of all ,you find the acceleration of the car, acceleration=(Vf-Vi)/t here, Vf=117*(5/18)m/sec=32.5m/sec then acceleration=(32.5-0)/0.977=33.265m/sec^2 now,applied force=mass*acceleration=945*33.265=31435.425N now, impulse=force*time=31435.425*.977=30712.41N-sec Hence change in linear momentum=impulse=30712.41N-sec=30712.41Kg.m/sec